A 10.0 mL sample of a 25% (m/v) KOH solution is diluted with water so that the final volume is 100.0 mL. Calculate the final concentration.
Answer #1

Answer : 2.5% final concentration

explanation :

We have been given, the initial volume of KOH solution Vi = 10 ml.

And, the initial concentration of the KOH solution (m/v) Mi= 25%.

since , according to the question, the KOH solution is diluted with water this means ,the more water is added to the original solution and that makes the original solution diluted and the concentration of solution decreases.

Again, the final volume of the solution after addition of water Vf = 100 ml.

Final concentration Mf = ?

To find out the final concentration we shall use the following formula-

MiVi = MfVf

10* (25%) = 100 ml *Mf

Mf = 10*25/100 %

Mf = 2.5%

Therefore, the final concentration of the KOH solution after addition of water will be 2.5% (m/v)

please upvote